If two inlet pipes are both open, they can fill a pool in 3 hours and 36 minutes. One of the pipes can fill the pool by itself in 6 hours. How long, in hours, would it take the other pipe to fill the pool by itself?
Answer by edjones(8007)

3 h 36 m = 3.6h
In one hr both pipes do 1/3.6 of the job.
in 1 hr the 1st pipe does 1/6 of the job.
1/6 + x = 1/3.6
x=1/3.6 - 1/6
=10/36 - 6/36 = 4/36 = 1/9 of the job done by the 2nd pipe in 1 hr.
9 hrs time it takes the 2nd pipe to fill the pool.
.
Ed